Y=2x+1, find the pmf of y ,E(y) and var(y)

Zhejiang University of Science and Technology 2. The joint pmf of X and Y is 2 6 0 4 0 Find the marginal distribution of Y and E(XY).
Answer #1

marginal distribution of y------

Y 1 2 3
P(Y=y) 1/6+1/6 1/6+0 1/4+1/4
P(Y=y) 1/3 1/6 1/2

E(XY) = ΣXY*P(x=x,Y=y)  

X Y P(X=x,Y=y) X*Y*P
0 1 1/6 0   
0 2 1/6 0   
0 3 1/4 0   
1 1 1/6 1/6
1 2 0    0   
1 3 1/4 3/4
total 11/12

so, E(XY) = 11/12 =0.92

----------------------------------------------------------------------------------------------------------------------------------------

for this question , Y=2x+1, find the pmf of y ,E(y) and var(y)

please define X